The critical path for the network activities shown below is __ with duration ___

Activity

Duration

Immediate Predecessors

A

2

-

B

4

-

C

1

A,B

D

6

A,B

E

3

B,C,D
Answer #1

The project network is shown below:

с A E B D

The project completion path is along A-C-E or B-D-E

Duration along path A-C-E = 2 + 1 + 3 = 6

Duration along B-D-E = 4 + 6 + 3 = 13

Since, longer time is along B-D-E,

The critical path for the network activities is along B-D-E with duration 13.